OpenStudy (naturesweetie):

Help with a couple questions please 1. The temperature of 100 g of liquid water in a calorimeter changes from 25°C to 50°C. How much heat was transferred? The specific heat of liquid water is 4.18 J/g-°C. 10.5 J 10.5 kJ 100 J 100 kJ 2. A 200 g block of a substance requires 520 J of heat to raise its temperature from 25°C to 45°C. Solve. 0.46 0.90 0.13 0.39

OpenStudy (naturesweetie):

Number two I got 0.13...

OpenStudy (aaronq):

use \(q=m*C*\Delta T\)

OpenStudy (naturesweetie):

So q= mass(100g) * C(4.18 J/g-C) * T(50-25=25C) This is for question one?

OpenStudy (aaronq):

yes

OpenStudy (naturesweetie):

I get when multiplying across 10,450... Then do I convert?

OpenStudy (naturesweetie):

the answer is 10.5 right? Is it kj or j? kj right?

OpenStudy (aaronq):

yes kJ

OpenStudy (naturesweetie):

Thank you very much!!